Radiometer Calibration Facility Log

 

 Date

Operator

Comments

20110826

Craig Webb

NREL called and let us know that enough data was collected to conclude borcal 2011-02.Will wait for

Calibration stickers before tearing down borcal 2011-02 instruments.

20110824

Craig Webb

Collected data today and weather is to be good tomorrow alsowill collect both days then let NREL

review the data.

20110712

Craig Webb

Havenít had weather for collecting borcal data, will be on the road for the next two weeksandno

Data will be collected.

20110706

Craig Webb

NREL stated that the data had a few problems with it so we will need another day or two of data.

20110630

Craig Webb

Started borcal 2011-02 today weather sunny but some haze through out the sky, could be from fires

Out west.Ran all day will let NREL check data.

20110629

Craig Webb

Did system cabling check and the light dark test for all the radiometers.

20110627

Craig Webb

Did inventory data sheet and inspection of radiometers for borcal 2011-02. Started installing

radiometers on tables and trackers.

20110614

Craig Webb

Borcal 2011-01 has been completed and the radiometers have been sent to the field.Will starting

to work on the next borcal.

20110516

Craig Webb

Fired up borcal andran most of the day started at 84 degrees and ran till 44 degree that afternoon,

Had to shut down due to clouds and wind.

20110510

Craig Webb

NREL called and stated they had the borcal system ready to go.Fired up the system and ran for

four hours till the clouds moved in and shut borcal down.

20110504

Craig Webb

Loaded the borcal 2011-01 data into the computer and completed the System light/dark test.Ran

a test session and a few problems were noted.Called NREL and they are looking into the system.

20110503

Craig Webb

Pulled radiometers and did the RCF Tracker layout and table layout for Borcal 2011-01.Started the

Radiometer Inventory and inspections.

20110418

Craig Webb

Did the data acquisition System functional Test and repaired 5 different locations.

20100809

Craig Webb

NREL responded to E-mail, they said that we should have enough data to conclude this borcal.

20100803

Craig Webb

Ran borcal today with both cavities with very few alarms.Ran for five hours till I had to shut

down cavity 30495, due to wind direction and speed.Started back up and running with one cavity

only till 83 degree that evening.Shut down due to farmerin field next to RCF.Let NREL know

that we ran all day long, should have enough data to conclude this borcal.

20100730

Craig Webb

Started borcal today and run for 8.5 hours, just after solar noon the clouds moved in and shut down borcal.

20100726

Craig Webb

Tried to run borcal but clouds started forming right after the first 30 minutes.

20100721

Craig Webb

Ran Borcal for about 45 minutes having a lot of errors and alarms,found the ventilator for

31101F3 to have a fan that was causing a heating problem and alarm condition.R/R ventilator

system up and running.

20100720

Craig Webb

Started borcal to do a test run and check the system out.Used one cavity only due to wind direction and speed.

20100714

Craig Webb

Input data into Borcal system to start Borcal 2010-02 and completed the light/dark test found one instrument that had the wrong location corrected.We are now waiting for the OK from NREL to start collecting data.

20100713

Craig Webb

Installed the rest of the radiometers for the current borcal.Double checked serial numbers and instruments leveling.Sent NREL a copy of Borcal 2010-02 table and tracker layout and inventory listing.

20100712

Craig Webb

Install all the NIPs and B/Ws for borcal 2010-02.

20100709

Craig Webb

Started getting instruments ready for Borcal 2010-02.

20100707

Craig Webb

NREL called and informed US that borcal 2010-01 was concluded at this time. Given instructions to tear down borcal 2010-01.

20100630

Craig Webb

NREL called and had us run a few hours and test a few instruments before concluding this borcal # 2010-01

20100605

Craig Webb

Ran borcal all day had some very good sessions and a few small clouds around solar noon, then sky cleared up and ran till 16:30est.With the total data collected will get with NREL to see if we enough data to to conclude this borcal.

20100530

Craig Webb

Ran for two hours before shutting down due to weather moving into the area.

20100529

Craig Webb

Ran till close to solar noon before clouds formed and shut down borcal.

20100528

Craig Webb

Ran for a few hours but clouds shut down borcal.

20100522

Craig Webb

Ran borcal today for about 10 hours today with a few clouds and alarms due to birds in the area.

20100512

Craig Webb

NREL called and they have fixed the problem with borcal 2010-01 and I installed the cavities and rean

I ran for two hours without any errors except for birds in the area.Waiting for good weather.

20100504

Craig Webb

Started taking data today but two of the STD radiometers were in a constant alarm and NREL had the borcal stopped till they had a chance to review the system data.

20100503

Craig Webb

NREL called and said I could edit the borcal system to reflect borcal 2020-01, finished and also completed the light dark test found one bad NIP, during the light/dark test.†† R/R nip and system checked out with out any problems.

20100426

Craig Webb

Started installing radiometers for borcal 2010-01 and finished installing them 4/28/10.

20100419

Craig Webb

Started radiometer inventory and inspection for borcal2010-01.

20100415

Craig Webb

Started mapping out Borcal tables & trackersfor borcal2010-01.

20100414

Craig Webb

Repaired the 6 bad positions found during the data Acquisition Functionaltest. Ready to start borcal inventory and inspection.

20100413

Craig Webb

Statedthe Data acquisition System Functional Test found 6 positions that need repaired.

20090823

Craig Webb

NREL called and we have enough data to conclude Borcal #200903.

20090821

Craig Webb

Ran borcal today and should have enough data to finish this borcal.

20090813

Craig Webb

Tried to finish borcal 2009-03 but clouds moved in and had to shut down.NREL called and stated that we only need early morning data to finish this borcal.

20090812

Craig Webb

Clouds cleared and started Borcal 2009-03 ran from 46 degrees to 83 degrees with no problem.

20090812

Craig Webb

Ran a test session today and everything is ready for Borcal Run.

20090810

Craig Webb

Radiometers installed on tables and the validation check done.Light dark test done also.

20090806

Craig Webb

Started setting up for Borcal 2009-03 to include the new radiometers for AMF #2

20090724

Craig Webb

NREL concluded Borcal 2009-02

20090723

Craig Webb

Ran Borcal today the data collected today might conclude this borcal will notify NREL and have them check the data quality.

20090716

Craig Webb

Ran borcal for a few hours.shortly after stating s bird shit on cavity 30495 and then data for that cavity started having data quality problems.Some bird shit got inside the cavity.Forwareded the info and a picture of it to NREL for their guidance on what to do with the cavity.

20090701

Craig Webb

Started running borcal 2009-02Ran for several hours but clouds moved in and shut down borcal.

20090629

Craig Webb

We are set up for borcal 2009-02.Instruments are installed and the light/dark test has been accomplished.Will run a test session this morning if the clouds part.

20090622

Craig Webb

Radiometers have been returned from the SGP sites and I will be working on getting Borcal 2009-02 ready for installation and layout.

20090603

Craig Webb

NREL called and said that Borcal 2009-01 was complete and they were sending the cal stickers our way and to start tearing down and getting ready for next Borcal.

20090519

Craig Webb

Ran Borcal today ran for three hours and wind shut down cavity 30495,ran with one cavity the rest of the day had good data sets and will let NREL review the data.

20090517

Craig Webb

Ran borcal all day today but had a few alarms and unstable conditions today.

20090421

Craig Webb

Ran for two hours today but clouds and contrails shut down the system.

20090420

Craig Webb

Ran borcal for two hours and clouds moved in and had to shut down.

20090414

Craig Webb

System test and verification of instruments.

20090402

Craig Webb

Finished installing all the radiometers for Borcal 2009-01 at this time, completed the light dark test.

20090330

Craig Webb

Craig and Rod starting installing radiometers on the tables and trackers at the RCF.

20090325

Craig Webb

Inventoried and inspected radiometers for borcal 2009-01.

20090324

Craig Webb

Did a mini test of the new borcal system.Seems to be working fine.

20090311

Craig Webb

NREL(Steve Wilcox @ Afsin) arrived on site and help with the new RCF Borcal Data system update.

20090218

Craig Webb

Received the new data loggers for the RCF Borcal system and started installing.

20090210

Craig Webb

Started to remove the old Borcal data logger system and install the new data interface boxes.

20080729

Craig Webb

Borcal 2008-02 Complete and radiometers selected to go out to their locations.

20080725

Craig Webb

NREL called and stated that the data collected looked OK, but donít tear down yet and to send the Helios to NREL for a quickpost calibration check since the PS had been changed out during the second borcal.

20080722

Craig Webb

Weather clear and borcal up and running. A few clouds in the area but not affecting borcal at this time ran for 13 hours today and will have NREL check the data to see if they have enough data to conclude this borcal.

20080721

Craig Webb

Started borcal and ran for six hours before farmers started burning wheat fields and smoke shut down borcal.

20080716

Craig Webb

Received the PS module and test it, installed and did a system test and everything is working now we are waiting for good weather to collect data.

20080714

Craig Webb

Tried to run borcal this morning but having problem with the data logger, unable to communicate with it, called NREL trouble shot and found the power supply module was bad.They will send a new module.

20080707

Craig Webb

Ran the light/dark test and run a test session to check out the borcal system.

20080701

Craig Webb

Started setting up for borcal 2008-02.Once all radiometers are installing we will do the light/dark test.

20080630

Craig Webb

Field Techs are changing out radiometers this week and when they return with the radiometers we will start the inventory and radiometer layout for Borcal 2008-02

20080609

Craig Webb

NREL called and the data is good enough to conclude Borcal 2008-01, Will start deciding where radiometer will be sent to for installation.

20080606

Craig Webb

Running both cavities data looks good ran for nine hours, Will talk with NREL to see if we have enough data to conclude this borcal.

20080604

Craig Webb

Started running Borcal after fog lifted sun at 46 degree, run one cavity due to wind, shut down at 1800 GMT due to clouds

20080603

Craig Webb

Weather broke and stated borcal running with one cavity only due to wind.Clouds moved in and shut down borcal at 1635 GMT.

20080525

Craig Webb

Weather has set in and no Borcal is possible.

20080516

Craig Webb

Clear morning ran for about four hours till cloud moved in and shut down borcal.

20080511

Craig Webb

Ran for 12 hours but data is not the best do to light haze though out the sky.

20080508

Craig Webb

Working with NREL we got both cavities reading like they should, will be ready to run borcal when weather breaks.

20080505

Craig Webb

Tried to run a test session but cavities were reading too low of an output.Working on the problem.Weather moved in and stopped the troubleshooting.

20080501

Craig Webb

NREL fixed the problem completed the light /dark test. Checked trackers for time and date stamp and adjusted.

20080430

Craig Webb

Finished installing radiometers and tried to does the light/dark test but program would not load up on the borcal computer.Notified NREL of problem.

20080429

Craig Webb

Finished inventory and started installing radiometers on trackers and tables.

20080428

Craig Webb

Helios returned from NREL and installed, Started system cable testing and found 9 bad cables.Repaired cables and retested.Started Borcal 2008-01 table lay-out and inventory.

20070820

Craig Webb

Just received word form NREL that, They have enough data to end Borcal 2007-02

20070813

Craig Webb

Ran Borcal today good data till about solar noon then shut down for two hours talk to NREL and they run the rest of the day and that would look over all the data collected on Tuesday morning and let us know if we have enough good data to conclude this borcal.

20070812

Craig Webb

Ran all day data in morning not to good but afternoon was good data.

20070811

Craig Webb

Ran all day data collected was good except around solar noon.

20070810

Craig Webb

Ran Borcal very good till about solar noon then small clouds were present for about two hours then cleared and afternoon was good data.

20070807

Craig Webb

Sky cleared around 1150 CST ran for 7 hours, had problem with cable going to 31121E6 had to repair cable.

20070806

Craig Webb

Ran Borcal for 9 hours had good data till field south of site started burning and shut borcal down.

20070725

Craig Webb

Finished working with NREL to get computer ready for borcal 2007-02, Completed the light-dark test and everything checked out.Ready for good weather so we can collect data.

20070723

Craig Webb

Finished installing the radiometers for borcal 2007-02.

20070718

Craig Webb

Started installing radiometers for Borcal 2007-02.

20070717

Craig Webb

Started the Borcal inventory log and inspecting the returned radiometers.

20070711

Craig Webb

Started laying out borcal 2007-02 and completed tracker and table lay-out.

20070702

Craig Webb

First set of radiometers for this yearís swap-out are ready for the field techs. To put them out in the field and return radiometers for the second borcal.

20070612

Craig Webb

NREL stated that there was enough data to conclude borcal 2007-01

20070608

Craig Webb

Started borcal and ran all day long, 13 hours.Data seem to be what NREL is looking for will check with NREL to see if we have enough data to conclude this borcal.

20070605

Craig Webb

Started borcal but as the sun started coming up clouds moved in and shut borcal down.

20070516

Craig Webb

Ran for two and a half hour then clouds moved in to shut down borcal.

20070510

Craig Webb

Test run #2 for this borcal, everything checks well.

20070504

Craig Webb

Started borcal today and ran a test session for 45 minutes, found that the humidity sensor was not working so the probe was replaced and as of this date we are ready to start Borcal 2007-01when weather permits.

20070502

Craig Webb

NREL Oked Configuration and a light dark test were completed, with satisfactory results.

20070430

Craig Webb

Loaded data into borcal computer for borcal 2007-01.Did a Configuration Audit so all data could be double checked for correctness.

20070423

Craig Webb

Finished installing radiometers on the roof top tables and trackers.

20070419

Craig Webb

Finished inspecting radiometers and started to install them on the RCF tables.

20060418

Craig Webb

Pulled radiometers off the shelf and started inspecting them and recording condition of each radiometer.Installed the NIPs on the RCF trackers today.

20070417

Craig Webb

Started working on Borcal 2007-01. Did table and tracker layout, sent a copy to NREL for approval.

20070416

Craig Webb

With the help of Rod Soper we conducted the system checkout and also conducted the cable test, found 8 bad locations after repairing these cable ran a retest of these locations and they all checked good.

20060829

Craig Webb

Borcal 2006-02 concluded at this time due to cavity controllers being returned to NREL for calibration.

20060827

Craig Webb

NREL called and requested that I send the cavity controllers to NREL to be calibrated before the up-coming NPC2006, therefore borcal will end this coming Friday if weather does not corporate.

20060822

Craig Webb

Weather still no cooperating therefore no borcal data taken.

20060815

Craig Webb

Didnít have conditions suitable for borcal.

20060812

Craig Webb

Ran for 7 hours, had good data but clouds shut borcal down one hour before solar noon.

20060808

Craig Webb

Ran for 3 Ĺ hours and clouds moved in, had to shut down.

20060806

Craig Webb

Ran all day but had some clouds pass through during the day.Some data is very good other times it is marginal but could be used if weather does not improve between now and the end of borcal season

20060731

Craig Webb

Ran for 9 hours before shutting down due to clouds

20060730

Craig Webb

Ran Borcal for three hours before clouds moved in and shut down borcal.

20060723

Craig Webb

Came in ran for 9 hours didnít get that morning and the haze and clouds shut borcal down a seventy degrees that afternoon.What data collected was very good.

20060720

Craig Webb

Ran for seven hour to day had to shut down just before solar non.

20060718

Craig Webb

Ran for a few hour today but clouds moved in.

20060717

Craig Webb

Ran Borcal for 6 sessions to day but clouds formed and shut down borcal.

20060714

Craig Webb

Ran ten sessions trying to find and fix problem with the cavity shutter opening and closing on it own and causing HF cavity to have a very lo output.†† Talked with Steve Wilcox at NREL and this borcal will be accomplished with cavity 29222 only.

20060713

Craig Webb

Borcal 2006-02 has been configured and instrument has been installed.The light dark test was conducted and instrument location verified.Waiting for weather to run borcal.

20060705

Craig Webb

Radiometers are ready to be exchanged on the southern run.We will start selecting radiometers to start Borcal 2006-02 and running them thru the blackbody.

20060616

Craig Webb

NREL notified us that Borcal 2006-01 had enough data to conclude this borcal.Will start changing out radiometers on the filed tech. Southern run and the central facility.

20060614

Craig Webb

Ran borcal all day and data seems to be enough to conclude this borcal.Will check with NREL in the morning.

20060613

Craig Webb

Trouble shooting 30495-cavity problem; found that cable connected to back of controller had a bad connection.Cleaned connector and ran 2 hours of testing and no problem.Ran borcal for 6 hours and everything ran very good.

20060609

Craig Webb

Started borcal today ran for a few hours then cavity 30945 had a problem with a low direct output.Then clouds moved in and closed borcal for the day.

20060606

Craig Webb

Condition for borcal has not been acceptable to this date.

20060602

Craig Webb

Started borcal good sky conditions ran for 7 hours and small cirrus started to form overhead and forced borcal to shut down.

20060525

Craig Webb

Ran close to eight sessions before clouds moved in a shut down borcal down.

20060520

Craig Webb

Started borcal ran 3 sessions and sky conditions shut borcal down.

20060519

Craig Webb

Ran all day but watts per meter squared just reached 830 watts. Had quite a few alarms with the MAS instruments.

20060518

Craig Webb

Run 5 sessions before clouds moved in and shut down the Borcal, just before solar noon.

20060517

Craig Webb

Run borcal today, first 4 sessions ran just fine, and after the 5-session cavity 30495 started giving a bad calibration and M factor. Closed borcal down and restarted but system was reading bad sensitivity for cavity 30495, Finally deleted cavity 30495 and continued to run borcal till clouds moved in.

20060516

Craig Webb

Run the light-dark test for all radiometers and run a test session for borcal 2006-01 System seems to be running Ok at this time.

20060515

Craig Webb

Loaded radiometers into borcal data base and configured system for borcal 2006-01

20060511

Craig Webb

Repaired the bad channels on the Borcal tables and trackers and retested them.Check OK.Install rest of the radiometers for this borcal.

20060510

Craig Webb

NREL has new version of Borcal ready for a test.Ran system test and checked all cables and connectors during system test.A few of the channels had high resistance and needed the cables repaired.

20060504

Craig Webb

Started inventor/inspection for the 8-48 radiometers and PSP.NREL working on new version of the Borcal Program.

20060502

Craig Webb

Installed the NIPís and checked the Brusag tracker clocks and positioning.

20060423

Craig Webb

Finished layout for Borcal 2006-01 and started radiometer inspection and inventory.

20060417

Craig Webb

Started working on Borcal 2006-01 radiometer layout.

20050725

Craig Webb

Called NREL and they are checking data to see if there is enough to conclude this borcal.

20050721

Craig Webb

Ran borcal 2005-02 today from daylight till dark, there were several alarms during the day. 31099F3 seems to be causing a problem.Found the fan on 31099E3 not working r/r fan and seems to be working fine.

20050720

Craig Webb

Finished installing the radiometer and did the light dark test and validation test.

20050719

Craig Webb

Installed the NIPís on trackers and aligned.Started installing the PSP

20050711

Craig Webb

Sorted the radiometers and did table mapping and tracker placement for the next borcal (2005-02) Completed the radiometer inventory and inspection data sheet.

20050705

Craig Webb

Took radiometers to shipping for NSA and TWP.Sent radiometers to the 10 sites in Kansas for the maintenance techs, to swap out this week.

20050620

Craig Webb

Ran all day today and we might have enough data for a borcal, NREL called and this borcal is finished

20050619

Craig Webb

Ran for 9 hours today, had a late start due to haze and light cirrus through the area.

20050615

Craig Webb

Ran for seven hours today till weather conditions shut down borcal.

20050604

Craig Webb

Run Borcal for several hours today till clouds moved in and had to shut down.

20050613

Craig Webb

Ran Borcal for 6 hours today from 1930 gmt till 2445 gmt should be good data.

20050603

Craig Webb

Borcal modified to run with PIR data collection to help decease the uncertainty.

20050509

Craig Webb

No borcaling due to weather and a BCR that is changing the BORCAL process.

20050430

Craig Webb

Ran BORCAL from daylight till dark.I will notify NREL about the data taken and let them review data and see if enough data was taken to conclude this borcal.

20050427

Craig Webb

Ran BORCAL for several hours today, shut down at 1840 GMT due to clouds moving in.NIP 31343E6 has a noisy output, notified NREL and awaiting their guidance on what to do.

20050425

Craig Webb

There was no BORCAL activity due to sky and weather conditions in local area this past week-end.

20050421

Craig Webb

NREL call and everything looks Ok at this time.Started BORCAL 2005 Ė01 and run for one hour till clouds moved in. The NIPS that were acting up seem to be working just fine.

20050418

Craig Webb

Installed the radiometers on the RCF roof top tables and leveled.Started the light dark test but didnít get finished till 4/19/05.Found three NIPS that were reading incorrect.Will do indepth testing tomorrow.Notified NREL of results and they are looking over the Configuration Audit Reportat this time.

20050415

Craig Webb

Installed Normal Incident radiometers on the Brusa Trackers and connected cables.

20050412

Craig Webb

Starting working on laying out BORCAL 2005-01. Completed tracker layout and the radiometer table layout.Completed the Radiometer inventory and inspection.

20050411

Craig Webb

Run Data Acquisition System Functional Test and faxed the hard copy to NREL for their review.Found several cables that checked bad but the problem was moisture in the back shell of the connector. Cleaned and dried the connector to correct the problems.

20041207

Craig Webb

The PIR blackbody calibration system is back up and operating.

20041104

Craig Webb

The PIR blackbody calibration system is down due to a faulty air-compressor piston.Parts are on order and it will be about two weeks before parts arrive. There are 12 PIRs that have been run thru the chamber since NPC 2004.This info will be forwarded to NREL for review and new calibration stickers/data.Once stickers are received PIR will be changed out at 6 locations in KS.

20040910

Craig Webb

NREL called and wanted the voltmeters used for BORCAL and the Blackbody calibration system to be sent in for calibration.Voltmeters were removed and sent to NREL for calibration.Till farther notice there will not be any PIR calibrated till after NPC 2004.

20040909

Craig Webb

NREL declared there was enough good data to conclude BORCAL 2004-02.

20040902

Craig Webb

NREL has looked at the data for Borcal 2004-02 and they think that there is enough good data to conclude BORCAL 2004-02 at this time.

20040829

Rod Soper

Started running borcal at 0955 GMT and run till 1800 GMT before clouds moved in and forced borcal to shut down.

20040807

Craig Webb

Started borcal at 0626 GMT and run just past solar noon when clouds started forming, shut down at 1445 GMT.

20040826

Craig Webb

Started running borcal at 0630GMT and ran till 1435 GMT till clouds moved in shut down.

20040825

Craig Webb

Ran borcal from 1945 GMT to 2415 GMT and had to shut down due to clouds

20040818

Craig Web

Started running borcal this morning at 0640 gmt till 1420 gmt when clouds moved in and shut borcal 2004-02 down.

20040813

Craig Webb

Started Borcal this morning and ran for about one hour before clouds formed and caused a shut down.

20040804

Craig Webb

After contacting NREL, they review the Borcal data and requested another morning of data be collected before concluding this Borcal.

20040801

Craig Webb

Ran Borcal from 80 degrees to 80 degrees but the solar noon session was not usable.

20040731

Craig Webb

Came in and ran for three hour before the clouds moved in.

20040713

Craig Webb

Started Borcal again today but due to a heavy haze on the horizon didnít get started till 1300 gmt and ran till clouds moved in and heavy smoke due to burning wheat fields.

20040712

Craig Webb

Started borcal at 85.5 degrees and ran till 1900 gmt. Shut down due to clouds and burning fields.

20040711

Craig Webb

Run through solar noon but had to close down due to clouds moving in.

20040709

Craig Webb

No progress made on Borcal 2004-02 due to weather conditions.

20040706

Craig Webb

Did not have weather compatible to run Borcal 2004-02 through out the holiday weekend.

20040702

Craig Webb

Changed out the measurement Assurance Standard PSP and installed the additional radiometers that NREL wanted to run through borcal 2004-02.Did light/dark test and completed testing.Waiting for good weather to continue with this borcal.

20040623

Craig Webb

Talked with NREL and they have finished updating the borcal computer.Input data for borcal 2004-02 into the computer system, finished and will start light dark testing tomorrow.Trouble shot the temp/hum probe problem found that we have temp but not humidity reading on both probes.

20040622

Craig Webb

Finished RCF table layout for Borcal 2004-02, and radiometer inspection. Started installing radiometers for this borcal.

20040611

Craig Webb

Turned over 27 radiometers to shipping to send off to TWP.Waiting for calibration stickers before send radiometers to NSA.

20040607

Craig Webb

Radiometers and Sirs programs are ready to be deployed this week, for the sites in Oklahoma.

20040602

Craig Webb

Received approval from NREL that the data collected for Borcal 2004-01 was enough to conclude this borcal.Will start selecting instrument for sites in the southern part of the cart site and edit the SIRS programs for these sites.After the programs are edited I will start tearing down this borcal set-up.

20040514

Craig Webb

Sent the PIR blackbody chamber data to Reda at NREL today for all the PIRs that belong to the TWP and NSA.When I get the calibration stickers I will have them shipped out to their locations.

20040512

Craig Webb

NREL called and said that the data looks good at this time and that they would get back with a more in-depth report.

20040505

Craig Webb

Started Borcal again today and ran till dark, was a very good day for data collection.We have about 42 sessions of data for this borcal.Notified NREL to have them check the quality of data that has been collected.

20040504

Craig Webb

Run Borcal for about 4 hours till clouds moved in and had to shut down.

20040503

Craig Webb

Received the parts to fix the black body dry air source.Installed and started the black body calibration again.At about 07:30 CST the clouds cleared out and I start Borcal 2004-01 ran till 1930 CST and shut down. Very good day.

20040430

Craig Webb

Didnít run borcal the rest of the week due to weather conditions.

20040427

Craig Webb

Started running borcal 2004 today ran from sunrise to sunset, should be good data.Will contact NREL and have them look at the data.When I came in this morning the blackbody dry air source had no output.Order part to fix it but no PIR calibration will be accomplished till the dry air source is fixed.

20040426

Craig Webb

Completed the system light dark test and aligned the NIPS install on the four trackers.

20040423

Craig Webb

Completed loading radiometer data and preformed the system test using the system test box.Saved the data as System Test 2004 & System Test2004B.

20040420

Craig Webb

Finished installing all the radiometers for borcal 2004-01. There is no sun so we could not align the NIPS that were install on the trackers.Started inputting data into the RCF computer.

20040419

Craig Webb

Sent the PIR data to NREL for 9 PIR from North Slopes, that has been through the blackbody calibration chamber.

20040416

Craig Webb

Started installing radiometers on the RCF tables and leveling them.

20040414

Craig Webb

Received approval from NREL for Borcal 2004-01 layout.Finished the radiometer inventory and inspection of all the radiometers that will be installed for borcal 2004-01.

20040409

Craig Webb

Started working on the RCF Radiometer Table layout and radiometer inventory and inspection for Borcal 2004-01. Finished the layout and faxed a copy to NREL for their approval.

20040115

Craig Webb

Ordered new cable and connectors to make and replace the control and data cables that connects the HF Cavity to the cavity controller.

20031104

Craig Webb

Radiometers from BORCAL 2003-02 have been exchanged in the rest of the ARM Site to finish exchange all radiometers except for the PIRs.The PIR blackbody system is back up and running after repair of the Lt-50 chiller and the dry air generator.

20031017

Craig Webb

NREL has review Borcal 2003-02 data and issued the calibration stickers.The new calibration stickers have been installed and the radiometers are ready to be installed at the ARM/CART sites.

20030929

Craig Webb

No change in status of BORCAL 2003-02 due to the workload of NREL, due to NPC-2003 Inter-cavity comparison.

20030909

Craig Webb

NREL had me remove the cavity controllers for cavity 29222 and 30495 and ship them to NREL for calibration before the upcoming NPC 2003.The controller for the blackbody calibration system was also sent in for calibration at this time.

20030908

Craig Webb

Notified NREL about the amount of data that had been collected over the weekend and had them check the data quality and they think we have enough data to halt the BORCAL at this time but not to disassemble any systems yet.

20030906

Craig Webb

Had another good day for collecting BORCAL Data today, except around 1315 local the farmer started to chisel his farmland and due to the wind and dust I suspended and capped cavity 30495.

20030905

Craig Webb

Had a very good day for collecting BORCAL data today ran from sun-up to sunset.

20030902

Craig Webb

No data taken over the weekend due to weather.

20030826

Craig Webb

Collected data till just after solar noon them clouds moved in and BORCAL was shut down.

20030824

Craig Webb

Collected data for about 7 hours shut down just after solar noon due to clouds moving in.

20030816

Rod Soper

Collected data until about 1100 CDT before clouds moved in and shut the BORCAL down.

20030815

Craig Webb

Collected data for 5 sessions before clouds moved in and shut the BORCAL down.

20030808

Craig Webb

We received 7 new 8-48 radiometers and we installed them in this BORCAL.The current BORCAL 2003-02 is now configured and ready to start collecting data, both night time and BORCAL data.

20030806

Craig Webb

We got the CM22 from Bob McCoy installed and a new configuration report done along with the light/dark test.Anthony Bucholtz was installing his instruments also and should be done by the close of business today.The shaded and unshaded PIRs were installed and hooked up to the spare data logger and that new system seems to be working OK at this time.NOTE:there is no 8-48 or NIP installed at this time, due to a shortage of calibrated instruments.

20030805

Craig Webb

Installing the visiting instruments and awaiting instruction on how and where to install the PIR radiometer, which is being discussed at this time.

20030731

Craig Webb

Received the Data Acquisition System Functional Test and procedures and have completed the testing and all channels have checked out.After the functional test the light/dark for each radiometer was completed.

20030725

Craig Webb

Most of the Arm/SGP radiometers have been installed at the RCF and now we are waiting for the Data Acquisition System Functional Test procedures before doing anymore configuring for BORCAL 2003-02

20030718

Craig Webb

Started installing the radiometers that will be in BORCAL 2003-02.

20030707

Craig Webb

Radiometers ready to go out into the field and be installed by the field-techs.Will start configuring the next BORCAL.

20030702

Craig Webb

NREL called and we have enough data to conclude BORCAL 2003-01.When the new calibration stickers arrive we will break down the RCF set-up and get the radiometers ready to go into the field.

20030627

Craig Webb

Collected BORCAL data all day should have enough data to conclude this BORCAL.Notified Steve Wilcox at NREL and will let him review all the BORCAL data to see if there is enough data to conclude BORCAL 2003-01.

20030617

Craig Webb

Started collecting data for BORCAL ran for three sessions and had to shut down to a farmer burning his wheat stubble.The smoke and ash was drifting toward the RCF shut down and removed the cavities.

20030608

Craig Webb

Didnít start collecting data till 1015 CST due to haze on the horizon.Ran till 1745 CST and started getting a constant alarm condition due to haze and dust on the horizon.

20030606

Craig Webb

Started collecting BORCAL data again at 1315 CST, and run till 1840 CST.Dust and haze on horizon causing constant alarms, quite collecting data for the day.

20030528

Craig Webb

Weather for running BORCAL and collecting data.Started collecting data at 0630 CST and ran till just after solar noon.Clouds moved in and BORCAL shut down for the day.

20030505

Craig Webb

Radiometers have been installed at the RCF for BORCAL 2003-01 and a light dark test has been accomplished.I will try to collect data today as long as the weather holds out.Before this BORCAL season NREL updated the BORCAL process in the computer system.

20030218

Craig Webb

All of the PIR from TWP have been run thru the Blackbody calibration chamber and the data has been sent to NREL for review and processing new calibration stickers.

20020819

Craig Webb

NREL has reviewed the data for BORCAL 2002-02 and there is enough data to conclude BORCAL 2002-02.They are working on getting new cal stickers for the instrument that went thru BORCAL 2002-02

20020814

Craig Webb

A Very good day for BORCAL ran from early morning till sunset.This should conclude BORCAL 2002-02; I will contact NREL and have them check the data quality.

20020807

Craig Webb

The weather had not been suitable for BORCAL work the last couple of days.

20020807

Craig Webb

I have just finished running all the spare Pyrgeometers located at the RCF thru the blackbody calibration chamber and have shipped all the data to NREL so they can review the data and make the new cal stickers.

20020805

Craig Webb

We are collecting data for BORCAL 2002-02 but seem to be lacking the early morning data required to complete this BORCAL, The weather is not co-operating as we get a constant alarm in the morning because of the haze on the horizon.Will keep trying and collecting data.

20020718

Craig Webb

The next four Pyrgeometers that are going to the TWP have been run thru the blackbody calibration chamber and the data has been sent to NREL for them to review the data and print the new calibration stickers.

20020718

Craig Webb

Ran BORCAL 2002-02 was able to get three good sessions before the clouds moved in and shutdown the BORCAL for today.I was able to get the 55-45 degree zenith angle for the A.M.

20020711

Craig Webb

Weather has not been suitable for BORCAL work; there might be a good day Sunday, July 14 2002.

20020711

Craig Webb

Received the radiometer from C. Cornwall and start running them thru the Blackbody calibration chamber one at a time.

20020708

Craig Webb

Weather has not been suitable for BORCAL work.

20020703

Craig Webb

BORCAL 2002-02 is ready to start collecting data.All instruments have been installed and the light/dark test has been completed.

20020628

Craig Webb

Worked on mapping out BORCAL 2002-02 instrument placement and inspected all radiometers.Fax Tom Stoffel the results and the inventory sheets for his review.Will start installing the radiometers Monday morning 7/1/02.

20020626

Craig Webb

I installed six PIR on the trackers along with our reference PIR for an outdoor comparison as ask by NREL.All six PIR have been through the Blackbody test chamber.Started collecting data at 1520 CST.

20020624

Craig Webb

Received the new calibration tags for all the instruments that went through BORCAL 2002-01, Installed the tags and sent the instruments belonging to TWP to Darwin as instructed by Bill Porch.Attached the new tags and ed-logged the programs for all the radiometers going to the EF sites in Oklahoma.Field Tech will start swapping out radiometers on 6/26/02; there are enough radiometers to cover the entire Elfís in Oklahoma.

20020614

Craig Webb

NREL is working on the new cal factors for instruments in BORCAL 2002-01.

20020613

Craig Webb

Had to shut down the blackbody calibration chamber due to a power failure at the ARM/CART site today.Will start it back up when power is restored and the humidity is brought back down to around 20%.

20020601

Craig Webb

I started running the TWPís Pyrgeometers thru the blackbody calibration today.I have three of them on hand at this time.

20020601

Craig Webb

Ran BORCAL 2002-01 again today.This should insure enough data to conclude this BORCAL but will let Steve Wilcox at NREL make the final decision.There is three days of data.When I suspended cavity 30495 due to wind speed and started back up the first three session were not graphing group #11 so I stopped the process and restarted it and everything worked fine.

20020531

Craig Webb

We had clear skies today and run BORCAL from 06:31 till 19:29 was a very good day to take data.I will contact NREL to have them check and see if we have enough data to conclude this BORCAL 2002-01.May take one more day of data because of problems we had at the start of this BORCAL.

20020518

Craig Webb

Was able to run BORCAL today till just before Solar noon then clouds moved in and BORCAL was shut down.

20020514

Craig Webb

Today I finished running the second set of 12 Pyrgeometers thru the black body calibration chamber.I am in the process of transferring the data collected from the first set of 12 Pyrgeometers installed on the trackers to NREL so they can process the data and produce the calibration stickers.Then I will remove the first set of 12 radiometers and install the second set of 12 radiometers.

20020513

Craig Webb

Ran BORCAL today, It was an uneventful day ran till 1945 CST.

20020509

Craig Webb

Ran BORCAL today first session ran good then clouds moved in and I shut down till the clouds dissipated.Started back up a 0930CST, 2 hours later group # 3 had the signal level for 6 out of 10 instruments drop to half of their value.This caused us to shut down the BORCAL and try to eliminate the problem.The problem was in the Helios chassis, found a bad circuit card connection.To late to start BORCAL again today.

20020430

Craig Webb

We have run six more Pyrgeometers thru the blackbody chamber, and have six more waiting to go thru the blackbody chamber.The first 12 Pyrgeometers are still on the trackers collecting data.

20020428

Craig Webb

Ran BORCAL 2002-01 today, the weather was good till 1500 CST.I was able to get thru solar noon with out any problems.

20020416

Craig Webb

The new revision for the BORCAL has been installed by NREL and C. Webb and R. Soper have installed all of the instruments that will be going through BORCAL 2002-01.The light/dark test has been completed.NREL will be going over the changes to the BORCAL program the next day or two with SGP/RCF personnel.

20020416

Craig Webb

The first twelve Pyrgeometers have been through the blackbody chamber and installed on trackers at the RCF.They will be collecting data for the next 2 to 3 weeks them the data will be sent to NREL for review/quality control check.

20020328

Craig Webb

Pyrgeometers 31236F3 and 31195F3 have been calibrated in the blackbody and installed on the tracker outside along with 29144F3 to check the precision of the blackbody calibration system.

20020321

Craig Webb

Pyrgeometer Blackbody calibration system installed at the RCF Facility March 20, 2002, by Ibrahim Reda and Tom Stoffel of NREL with the help of Dan Nelson and Craig Webb.Training was conducted and the system turned over to RCF personnel for use.

20010910

Craig Webb

Another very good day for collecting data.After today I will contact NREL and have them check and see if there is enough data to conclude BORCAL 2001-02.

20010909

Craig Webb

Fired up BORCAL 2001-02 today and ran all day, collected some very good data.

20010829

Craig Webb

Just got thru the zenith angle of 55 Ė to Ė45 and clouds moved in and ended the BORCAL for the day.

20010828

Craig Webb

Fired up BORCAL 2001-02 was able to run till the end of solar noon then clouds moved in.

20010827

Craig Webb

Started BORCAL at 0610 and ran all day long.One very good day.

20010821

Craig Webb

Everything is ready for BORCAL 2001-02.Radiometers were installed and the database was created.The light dark test was completed and one test session was conducted.Just waiting on weather to proceed.

20010816

Craig Webb

All the radiometers that were calibrated have been installed in the field.The radiometers that were returned to the RCF will be inspected and installed on the RCF tables.

20010803

Craig Webb

Steve Wilcox called and informed me that there was enough data to conclude BORCAL 2001-01.There will be about a ten-day period between the next BORCAL # 2001-02.

20010723

Craig Webb

Good day for BORCAL.Ran the whole day.Notified Steve Wilcox to see if there is enough data to complete this BORCAL.

20010720

Craig Webb

Good morning for BORCAL, clouds caused to shut down just before solar noon.NREL installed revision 0.20 this morning.

20010719

Craig Webb

Run BORCAL 2001-01 till after solar noon when clouds moved in.

20010718

Craig Webb

Started BORCAL today after clouds lifted.At a zenith angle of 56, configured to run with one cavity due to wind speed and direction.Clouds caused shut down of BORCAL for the day.

20010712

Craig Webb

Conditions suitable for running BORCAL till 1345 CST.

20010711

Craig Webb

Conditions are fair for running BORCAL again today.Conditions were favorable until a zenith of 70 degrees in the pm.

20010710

Craig Webb

Second day of running for BORCAL 2001-01

20010709

Craig Webb

First day following RCC upgrade that was appropriate for BORCAL.

20010625

Dan Nelson

NREL (Steve Wilcox and Afshin Andreas) is on site to upgrade the RCC system.Work includes new software, a new PC, and new custom-built Active Cavity Radiometer controllers.The first BORCAL of 2001 is expected to begin immediately following the completion of these upgrades.

20000228

 Dan Nelson

 BORCAL season officially begins today.  However, NREL has discovered a Y2K issue with the RCC software that runs the calibration process.  Currently, they estimate that new software will be available by the end of April 2000.  Until that time, SGP is on hold with regard to radiometer calibration.

19990802

Dan Nelson

All BORCAL 1999-02 results sent to NREL (Steve Wilcox) for evaluation.

 19990730

 Dan Nelson

Conditions favorable for all day calibration.Diffuse component getting above 20% in the late afternoon and early evening halted BORCAL activities around a zenith of 70d.

 19990729

Dan Nelson

Conditions favorable for morning calibration.Few scattered thin cirrus moving off to the east.Building cumulus halted activities just before solar noon.

19990725

Dan Nelson

Conditions favorable for afternoon and early evening calibration.

19990724

Dan Nelson

Conditions favorable for afternoon and early evening calibration.

19990721

Dan Nelson

Conditions favorable for morning calibration.Building cumulus halted activities just before solar noon.

 19990720

 Dan Nelson

 Conditions favorable for morning calibration.Building cumulus and smoke from field burning to the SE halted activities just before solar noon.

19990719

Dan Nelson

Conditions favorable for morning calibration.Building cumulus and smoke from field burning to the SE halted activities just before solar noon.

19990718

Dan Nelson

Conditions favorable for morning calibration.Building cumulus halted activities just before solar noon.

19990713

Dan Nelson

Conditions nearly perfect for a full days calibration.Unfortunately, equipment problems have cost us several morning sessions, including 44-55.

19990709

 Dan Nelson

Conditions favorably for calibration through solar noon.Significant thin Cirrus halted work just pas solar noon.

19990708

 Dan Nelson

BORCAL 1999-02 began this morning.Conditions are predicted to be favorable for most of the day.

 19990629

Dan Nelson

SIRS PSPs and NIPs are being exchanged this week for the southern half of the SGP CART site.Additionally, problematic SIRS instruments at E09 & E10 are also being exchanged.SRS C1 (Test Bed) is also in the process of being exchanged.So far, we are on schedule to begin the next BORCAL (1999-02) near the end of next week.
 
 

Last week, Jim Treadwell (NREL) was on site to perform testing of the RCC system.Certain anomalies within the data for BORCAL 1999-01 prompted these tests.Jim found two loose connections on data connectors feeding into the Helios datalogger, which he corrected.Steve Wilcox, creator of the RCC program, suspects that these two problems coupled with our (Site Ops) improper removal of an instrument under test could yield spurious data.In the future, if an instrument (other than a cavity radiometer) fails during any phase of an ongoing BORCAL, it should be "Deactivated" rather than "Suspended" to avoid this problem.

 19990621

Dan Nelson

Received notification from NREL that BORCAL 1999-01 is complete and that the final report is ready for signature.Calibration tags should be ready this week.Radiometer exchange could begin next week and the next BORCAL started on or about 19990709.

 19990525

 Dan Nelson

 BORCAL results zipped up and shipped to Steve Wilcox at NREL for evaluation.Pending his review, BORCAL activities have been suspended.

19990524

Craig Webb

BORCAL started this morning at 0700CST.Cirrus drifted in within a couple of hours to halt BORCAL efforts.

19990519

Dan Nelson

BORCAL started this morning at 0730CST.  Little cirrus low to the east prevented an earlier start.

19990518

Dan Nelson

BORCAL started this morning at 0700CST.

19990515 thru 
19990517

Dan Nelson

No BORCAL today.  Total OVC.

19990514

Craig Webb

No BORCAL today.  Total OVC.

19990513

Dan Nelson

No BORCAL today.  Significant cirrus overhead.

19990512

Dan Nelson

BORCAL system brought back on line after restoring portions of the RCC PC hard drive from tape backup.  Calibration from just prior to solar noon until mid afternoon was possible before significant cumulus developed.

19990507

Dan Nelson

BORCAL system failed this morning.  There appears to be some sort of corruption within the RCC program.  NREL has been called to address this issue.

19990501 thru 
19990506

Dan Nelson

No BORCAL.  Weather conditions unfavorable for operation.

19990415 thru 
19990530

Dan Nelson

BORCAL when possible.  I will be away on vacation.  Craig Webb will be responsible for operating the RCF in my absence.  Craig does not have access to edit this page at the present.  When I return from vacation, I will arrange for him to have access to edit this page.

19990405

Dan Nelson

Setup continues for SGP BORCAL 1999-01.  Pre calibration testing will take place later this week.

19990330

Dan Nelson

Tom Stoffel (NREL) gave week's STM for Site Operations to prepare for the 1999 SGP BORCAL season approval at last.  This work will require several days and should finish up sometime next week.

19980914

Dan Nelson

With the end of the calibration season nearly upon us and given that any radiometers calibrated during this BORCAL season would require calibration again next spring, Tom Stoffel and I have elected to forego further calibrations this year.  Look for new entries to this log early next spring.

19980826

Dan Nelson

BORCAL 1998-02 complete!  The following is an excerpt from a message sent to us by Steve Wilcox of NREL following his review of the RCC data we recently sent to them. 

Dan, 

We've looked at the BORCAL.  Looks good!  The results for the MAS and control instruments are consistent with previous BORCALs.  There is a gap of data in the afternoon, but it appears that each afternoon bin on either side has some data, so the effect is minimal. 

Questions: 

What happened to the humidity sensor?  It looks like it failed part way through. 

We would like a copy of all session logs (copy and mail to us; 
conventional mail is okay). 

Go ahead and tear down.  We'll prepare the calibration report here and make copies for distribution.  We'll also produce the calibration stickers here for each instrument and send them to you. 

Thanks for all the good work. 

Steve

19980825

Dan Nelson

No BORCAL.  Scattered low cumulus.  Some morning high cirrus.

19980824

Dan Nelson

PM BORCAL attempt.  Data collected to this point sent to NREL for review.

19980823

Dan Nelson

No BORCAL.  Scattered low cumulus.

19980822

Dan Nelson

No BORCAL.  Scattered low cumulus.

19980821

Dan Nelson

No BORCAL.  Scattered low cumulus.

19980820

Dan Nelson

No BORCAL.  Scattered low cumulus.

19980819

Dan Nelson

With adequate AM BORCAL samples in hand, we will now concentrate on obtaining good PM data.  Up to this point, our attempts to gather PM data have been foiled by low cumulus forming around solar noon.  The NWS is calling for mostly sunny for the entire day today.  With any luck, we may just get our first PM data in a good many days.

19980818

Dan Nelson

AM BORCAL attempt.  Clear skies this morning.  PM remains questionable at this point, but we will run RCC as long as the skies cooperate. 

Note:  Some PM BORCAL data collected.  Few scattered low cumulus from time to time.

19980817

Dan Nelson

AM BORCAL attempt.  Clear skies this morning.  PM remains questionable at this point, but we will run RCC as long as the skies cooperate. 

Note:  Afternoon clouds canceled BORCAL.

19980816

Dan Nelson

AM BORCAL attempt.  Clear skies this morning.  PM remains questionable at this point, but we will run RCC as long as the skies cooperate. 

Note:  Afternoon clouds canceled BORCAL.

19980815

Dan Nelson

AM BORCAL attempt.  Clear skies this morning.  PM remains questionable at this point, but we will run RCC as long as the skies cooperate. 

Note:  Afternoon clouds canceled BORCAL.

19980814

Dan Nelson

AM BORCAL attempt.  Clear skies this morning.  PM remains questionable at this point, but we will run RCC as long as the skies cooperate. 

Note:  Afternoon clouds canceled BORCAL.

19980813

Dan Nelson

No BORCAL.  Significant cloud cover.

19980812

Dan Nelson

No BORCAL.  Significant thin cirrus.

19980811

Dan Nelson

No BORCAL.  Significant cloud cover.

19980808 thru 
19980810

Dan Nelson

No BORCAL.  Significant cloud cover.

19980807

Dan Nelson

AM BORCAL attempt.  Clear skies this morning.  PM remains questionable at this point, but we will run RCC as long as the skies cooperate. 

Note:  Afternoon clouds canceled BORCAL.

19980806

Dan Nelson

AM BORCAL attempt.  Clear skies this morning.  PM remains questionable at this point, but we will run RCC as long as the skies cooperate. 

Note:  Afternoon clouds canceled BORCAL.

19980805

Dan Nelson

No BORCAL.  Significant cloud cover.

19980804

Dan Nelson

No BORCAL.  Significant thin cloud cover.

19980803

Dan Nelson

No BORCAL.  Significant cloud cover.

19980731

Dan Nelson

No BORCAL today.  Total cloud cover with intermittent showers. 

Based on the attached NWS forecast for Grant county, conditions will also not be favorable over the weekend for BORCAL activities. 

ALFALFA OK-BLAINE OK-GARFIELD OK-GRANT OK-KAY OK-KINGFISHER OK- 
LOGAN OK-MAJOR OK-NOBLE OK-PAYNE OK- 
INCLUDING THE CITIES OF...ENID OK...GUTHRIE OK...PONCA CITY OK... 
STILLWATER OK 
430 AM CDT FRI JUL 31 1998 

.TODAY...MOSTLY CLOUDY WITH A 30 PERCENT CHANCE OF THUNDERSTORMS. 
HIGH AROUND 90.EAST WIND 10 TO 20 MPH. 
.TONIGHT...MOSTLY CLOUDY WITH A 40 PERCENT CHANCE OF THUNDERSTORMS. 
LOW NEAR 70.EAST WINDS AROUND 10 MPH. 
.SATURDAY...PARTLY CLOUDY WITH A 30 PERCENT CHANCE OF THUNDERSTORMS. 
HIGH IN THE MID 90S.SOUTHEAST WIND 10 TO 15 MPH. 
.SATURDAY NIGHT...PARTLY CLOUDY WITH A 20 PERCENT CHANCE OF 
THUNDERSTORMS.LOW IN THE LOWER 70S. 

.EXTENDED FORECAST... 
.SUNDAY...PARTLY CLOUDY.HIGH IN THE UPPER 90S. 
.MONDAY...MOSTLY CLEAR.LOW IN THE LOWER 70S AND HIGH 100 TO 105. 
.TUESDAY...PARTLY CLOUDY WITH A SLIGHT CHANCE OF THUNDERSTORMS.LOW 
IN THE LOWER 70S AND HIGH IN THE UPPER 90S.

19980730

Dan Nelson

No BORCAL.  Significant cloud cover.

19980729

Dan Nelson

No BORCAL.  Total cloud cover with intermittent showers.

19980728

Dan Nelson

No BORCAL.  Total cloud cover with intermittent showers.

19980723- 
19980727

Dan Nelson

No BORCAL.  Significant cloud cover.

19980722

Dan Nelson

No BORCAL today.  Significant clouds to the north and west.

19980721

Dan Nelson

Late AM BORCAL today.PM BORCAL halted by rapidly forming cumulus.

19980720

Dan Nelson

AM BORCAL good today.  Parts of PM BORCAL questionable.

19980719

Dan Nelson

AM BORCAL good today.  Parts of PM BORCAL questionable.

19980718

Dan Nelson

AM BORCAL good today.  Parts of PM BORCAL questionable.

19980717

Dan Nelson

No BORCAL today.  Mid and upper level clouds are currently present and are predicted to persist throughout the day.

19980716

Dan Nelson

No BORCAL today.  Mid and upper level clouds are currently present and are predicted to persist throughout the day.

19980715

Dan Nelson

AM BORCAL was successful up to just prior to solar noon.  Rapidly forming low-level cumulus halted our activities.

19980714

Dan Nelson

AM BORCAL attempted and soon aborted when significant cumulus formed.  Surface humidity is so high right now that clear mornings rapidly degrade.

19980713

Dan Nelson

No BORCAL today.  Nor does the NWS forecast for the remainder of the week appear favorable for BORCAL activities. 

GARFIELD OK-GRANT OK-KAY OK-NOBLE OK- 
535 AM CDT MON JUL 13 1998 

. TODAY...PARTLY CLOUDY WITH A 20 PERCENT CHANCE OF SHOWERS AND 
THUNDERSTORMS. HIGH IN THE MID 90S. VARIABLE WIND 5 TO 10 
MPH...BECOMING NORTHEAST AT 10 TO 15 MPH THIS AFTERNOON. 
.TONIGHT...PARTLY CLOUDY WITH A 20 PERCENT CHANCE OF SHOWERS AND 
THUNDERSTORMS. LOW IN THE MID 70S. EAST WIND 5 TO 10 MPH. 
.TUESDAY...PARTLY CLOUDY. HIGH IN THE MID 90S. EAST WIND 10 TO 15 
MPH. 
.TUESDAY NIGHT...PARTLY CLOUDY WITH A 20 PERCENT CHANCE OF SHOWERS 
AND THUNDERSTORMS. LOW IN THE LOWER 70S. 

.EXTENDED FORECAST... 
.WEDNESDAY AND THURSDAY...PARTLY CLOUDY WITH A SLIGHT CHANCE OF 
SHOWERS AND THUNDERSTORMS. LOW IN THE LOWER 70S AND HIGH IN THE UPPER 
90S. 
.FRIDAY...MOSTLY CLEAR. LOW IN THE LOWER 70S AND HIGH NEAR 100.

19980710

Dan Nelson

Shade & unshade testing prior to commencement of SGP BORCAL 1998-02 was performed today.  All systems are go for starting this activity and all we require now are clear skies. 

Today will not be a clear sky day at the SGP RCF.  Also, NWS predictions for clear skies over the weekend do not appear favorable for BORCAL activities.  Therefore no weekend calibration work will be scheduled.

19980709

Dan Nelson

Preparations for SGP BORCAL 1998-02 continue this week.  The start of 1998-02 could begin by the week's end.  However, based on the attached NWS zone forecast, the remainder of the week appear to be questionable for BORCAL activities.  The weekend looks good right now, but we will continue to monitor the NWS forecast for the remainder of the week to see if this prediction holds. 

 ALFALFA OK-GARFIELD OK-GRANT OK-KAY OK-MAJOR OK-NOBLE OK- 
430 AM CDT THU JUL 9 1998 

.TODAY...PARTLY CLOUDY WITH A 20 PERCENT CHANCE OF SHOWERS AND 
THUNDERSTORMS. HIGH IN THE UPPER 90S. SOUTH TO SOUTHEAST WIND 5 TO 
15 MPH. 
.TONIGHT...PARTLY CLOUDY WITH A 30 PERCENT CHANCE OF THUNDERSTORMS. 
LOW IN THE MID 70S. SOUTH WIND AROUND 10 MPH. 
.FRIDAY...PARTLY CLOUDY WITH A 20 PERCENT CHANCE OF SHOWERS AND 
THUNDERSTORMS. HIGH IN THE UPPER 90S. SOUTH WIND AROUND 10 MPH. 
.FRIDAY NIGHT...PARTLY CLOUDY WITH A 30 PERCENT CHANCE OF SHOWERS AND 
THUNDERSTORMS. LOW IN THE MID 70S. 

.EXTENDED FORECAST... 
.SATURDAY...MOSTLY CLEAR. HIGH IN THE UPPER 90S. 
.SUNDAY...MOSTLY CLEAR. LOW IN THE MID 70S AND HIGH 100 TO 105. 
.MONDAY...MOSTLY CLEAR. LOW IN THE MID 70S AND HIGH NEAR 100.

19980708

Dan Nelson

Preparations for SGP BORCAL 1998-02 continue this week.  The start of 1998-02 could begin by the week's end. 

NREL has requested that we exchange our Measurement Assurance Standards (MAS), returning the set used in SGP BORCAL 1998-01 to SSRL for calibration in their upcoming BORCAL.  To that end, we have returned MAS set #2 and installed MAS set #1 for the upcoming BORCAL.

19980707

Dan Nelson

Preparations for SGP BORCAL 1998-02 continue this week.  The start of 1998-02 could begin by the week's end.

19980706

Dan Nelson

Preparations for SGP BORCAL 1998-02 continue this week.  The start of 1998-02 could begin by the week's end.

19980702

Dan Nelson

Preparations for SGP BORCAL 1998-02 began this morning with inventory and inspection of all recently exchanged NIP and PSP radiometers.  This work will continue for several days and include RCC channel and deck location assignment.

19980629

Dan Nelson

Radiometer calibration tags and reports for SGP BORCAL 1998-01 were received from NREL last week.  These tags were installed and the radiometers were assigned for field exchange to take place at all southern EF SIRS during this week.  Additionally, we will also exchange radiometers at the four northern EF SIRS and the CF (E13) during this week. 

Following the 4th of July holiday, we will begin preparations for SGP BORCAL 1998-02.  This work should be completed by 19980710 and data gathering will commence at our first available opportunity. 

19980603

Dan Nelson

BORCAL 1998-01 data gathering is complete. NREL is satisfied with the data gathered thus far. 

Site Operations will generate and forward the final BORCAL 1998-01 report to NREL for processing.  NREL will prepare calibration tags and forward them to the site for installation onto the appropriate radiometers.  Following receipt and installation of the calibration tags (2-3 weeks from today), Site Operations will assign the freshly calibrated radiometers so they may be exchanged for 50% of the existing fielded radiometers.  These returned instruments will be prepared for the next BORCAL. 

Expect no further entries to this log until the RCF operators begin preparations for BORCAL 1998-02 sometime in July.

19980529 to 
19980602

Dan Nelson

BORCAL 1998-01 data forwarded to NREL for evaluation. 

There will be no further data gathering until these results are analyzed.

19980528

Dan Nelson

PM BORCAL.  Uniform Haze, but cloudless.

19980527

Dan Nelson

No BORCAL.  Unfavorable atmospheric conditions.

19980519 to 19980526

Dan Nelson

No BORCAL.  Unfavorable atmospheric conditions.

19980518

Dan Nelson

No BORCAL.  Partial OVC and moderate haze present.

19980517

Dan Nelson

No AM BORCAL.  Significant clear sky haze present.  PM BORCAL is also questionable.

19980516

Dan Nelson

BORCAL AM & PM.  Severely clear skies potentially offers us one of the best calibration days of the season so far. 

BORCAL update.  Murphy's law of solar radiometer calibration has struck us down.  Sometime just after solar noon, solar tracker #1 began acting up.  Azimuthal rotation was not smooth in either clock or sun modes.  The tracker would lag behind by several degrees and then slowly catch up.  This solar tracker points our windowless HF-ACR and both of the diffuse reference radiometer's shading assembly.  Attempting BORCAL with this type of solar tracking error condition will lead to significantly erroneous calibration results. 

The SGP RCF maintains a spare Brusag solar tracker on hand (good foresight on NREL's part), so I elected to exchange the spare for the faulty unit.  This effort required most of the afternoon because of the significant hardware exchange and precise alignment requirements associated with this activity.  The new unit was brought online at about 75 degrees zenith angle.  Both clock and solar tracking modes yielded pointing values within <0.2 degrees of one another, indicating extremely close initial alignment.  The tracker was placed in solar mode as recommended by the manufacturer and left for the night.  Attempting further calibration at this late solar hour was deemed not worth the bother.  I will check the solar pointing accuracy of this tracker in the morning.

19980515

Dan Nelson

PM BORCAL will be attempted despite high southerly winds.  NREL has outlined a method for suspending the windowless HF-ACR from the BORCAL during these frequent high wind, but very clear sky days.  The HF-ACR can be reinstated into the ongoing BORCAL when wind conditions allow; typically <6-7 m/s from the southerly direction.

19980514

Dan Nelson

No BORCAL.  Completely OVC.

19980513

Dan Nelson

No BORCAL.  Southerly winds in excess of 8 m/s preclude windowless HF-ACR stability, raising acceptable error parameters beyond allowable limits.

19980512

Dan Nelson

AM BORCAL.  The NWS forecast below appears not to reflect the actual local weather conditions.

19980511

Dan Nelson

AM BORCAL attempted, but strong southerly winds are causing significant windowless ACR drift. 

The attached NWS zone forecast shows the remainder of the week to be questionable for BORCAL. 

GARFIELD OK-GRANT OK-KAY OK-NOBLE OK- 
1030 AM CDT MON MAY 11 1998 

.TODAY...PARTLY CLOUDY AND BECOMING WINDY. HIGH IN THE MID TO UPPER 
80S. SOUTH WIND INCREASING TO 20 TO 30 MPH AND GUSTY. CAUTION ADVISED 
ON AREA LAKES. 
.TONIGHT...PARTLY CLOUDY WITH A 20 PERCENT CHANCE OF SHOWERS AND 
THUNDERSTORMS. LOW NEAR 60. SOUTHEAST WIND 10 TO 15 MPH. 
.TUESDAY...PARTLY CLOUDY. HIGH IN THE MID 80S. SOUTH WIND 5 TO 10 MPH 
SHIFTING TO THE NORTH IN THE MORNING. 
.TUESDAY NIGHT...PARTLY CLOUDY WITH A 20 PERCENT CHANCE OF 
THUNDERSTORMS. LOW IN THE LOWER 60S. 

.EXTENDED FORECAST... 
.WEDNESDAY...PARTLY CLOUDY. HIGH IN THE UPPER 80S. 
.THURSDAY...PARTLY CLOUDY WITH A SLIGHT CHANCE OF THUNDERSTORMS. LOW 
IN THE LOWER 60S AND HIGH IN THE UPPER 80S. 
.FRIDAY...PARTLY CLOUDY WITH A CHANCE OF THUNDERSTORMS. LOW IN THE 
LOWER 60S AND HIGH IN THE MID 80S.

19980510

Dan Nelson

BORCAL attempted, but afternoon clouds halted our work.

19980509

Dan Nelson

No BORCAL.  Significant Cumulus and scattered showers.

19980508

Dan Nelson

No BORCAL.  Significant Cumulus and scattered showers.

19980507

Dan Nelson

No AM BORCAL.  Significant Cirrus building to the south and west. 

PM BORCAL looks promising.  RCF is within a clear wedge of air between two large weather system to the north and south of us.

19980506

Dan Nelson

No BORCAL.  Presently overcast with same predicted for the remainder of the day.

19980505

Dan Nelson

No BORCAL.  Presently overcast with same predicted for the remainder of the day.

19980504

Dan Nelson

PM BORCAL is possible.  Clouds have moved out of the local area.  Satellite imagery and radar loop indicate that we should be in clear sky conditions for most of the afternoon. 

No AM BORCAL based on the attached NWS zone forecast.  Chance of BORCAL for the remainder of the week also questionable at this point based on the same forecast. 

GARFIELD OK-GRANT OK-KAY OK-KINGFISHER OK-LOGAN OK-NOBLE OK-PAYNE OK- 
430 AM CDT MON MAY 4 1998 

.TODAY...PARTLY CLOUDY. HIGH NEAR 80. SOUTH WIND 5 TO 15 MPH. 
.TONIGHT...PARTLY CLOUDY WITH A 30 PERCENT CHANCE OF SHOWERS AND 
THUNDERSTORMS. LOW IN THE MID 50S. SOUTHEAST WIND 10 TO 15 MPH 
SHIFTING TO THE NORTHEAST TOWARD MORNING. 
.TUESDAY...MOSTLY CLOUDY WITH A 20 PERCENT CHANCE OF SHOWERS AND 
THUNDERSTORMS. HIGH IN THE LOWER 70S. NORTHEAST WIND 10 TO 15 MPH. 
.TUESDAY NIGHT...PARTLY CLOUDY WITH A 20 PERCENT CHANCE OF SHOWERS 
AND THUNDERSTORMS. LOW IN THE LOWER 50S. 

.EXTENDED FORECAST... 
.WEDNESDAY THROUGH FRIDAY...PARTLY CLOUDY WITH A CHANCE OF SHOWERS 
AND THUNDERSTORMS. HIGH IN THE LOWER AND MID 70S. LOW IN THE UPPER 
50S.

19980503

Dan Nelson

AM BORCAL today.  PM BORCAL somewhat dependent on local winds and advancing clouds. 

PM BORCAL successful through about 1600 CDST.

19980501

Dan Nelson

BORCAL attempt based on the attached NWS zone forecast.  Fog will obviously delay start of BORCAL. 

GARFIELD OK-GRANT OK-KAY OK-NOBLE OK-PAYNE OK- 
430 AM CDT FRI MAY 1 1998 

.TODAY...FOG EARLY THIS MORNING...THEN MOSTLY SUNNY. HIGH IN THE MID 
70S. LIGHT WEST WIND.

19980430

Dan Nelson

No BORCAL.  Broken to scattered Cumulus all day.

19980429

Dan Nelson

No BORCAL.  The forecast calls for rain and continued cloudiness through 19980430.

19980428

Dan Nelson

No BORCAL.  The forecast calls for rain and continued cloudiness through 19980430.

19980427

Dan Nelson

No BORCAL.  The forecast calls for rain and continued cloudiness through 19980429.

19980424

Dan Nelson

No BORCAL.  Moderate Cumulus and gusting winds to 10 m/s. Additionally, based on the attached IWIN forecast, plans for weekend BORCAL have been scrapped. 

ALFALFA OK-GARFIELD OK-GRANT OK-KAY OK-MAJOR OK-NOBLE OK- 
430 AM CDT FRI APR 24 1998 

.TODAY...MOSTLY SUNNY AND WINDY. HIGH IN THE LOWER 80S. SOUTH WIND 15 
TO 25 MPH AND GUSTY. CAUTION IS ADVISED ON AREA LAKES. 
.TONIGHT...FAIR SKIES AND WINDY. LOW IN THE LOWER 60S. SOUTH WIND 15 
TO 25 MPH. 
.SATURDAY...PARTLY CLOUDY AND WINDY WITH A 20 PERCENT CHANCE OF 
THUNDERSTORMS. HIGH IN THE MID 80S. SOUTH WIND 15 TO 25 MPH AND 
GUSTY. 
.SATURDAY NIGHT...MOSTLY CLOUDY WITH A 40 PERCENT CHANCE OF 
THUNDERSTORMS. LOW NEAR 60. 

.EXTENDED FORECAST... 
.SUNDAY...THUNDERSTORMS LIKELY. HIGH IN THE UPPER 60S. 
.MONDAY...MOSTLY CLOUDY WITH A CHANCE OF SHOWERS AND THUNDERSTORMS. 
LOW IN THE UPPER 40S AND HIGH NEAR 70. 
.TUESDAY...PARTLY CLOUDY. LOW IN THE MID 40S AND HIGH IN THE LOWER 
70S.

19980423

Dan Nelson

AM BORCAL.  No clouds.  Winds intermittent at 0-2 m/s. 

PM BORCAL scrapped to determine the cause of an RCC software problem.  RCC was improperly applying a window correction factor to the windowless HF Absolute Cavity radiometer (30495).  With the problem fixed, we can restart BORCAL activities at our next clear opportunity.  Some question of data validity exists for the period when this problem was occurring; essentially all BORCAL 1998-01 up to this point.

19980422

Dan Nelson

AM BORCAL.  No clouds.  Winds intermittent at 0-2 m/s. 

No PM BORCAL.  Scattered clouds drifting in and winds increasing to about 6-7 m/s.  See the attached IWIN forecast below. 

GARFIELD OK-GRANT OK-KAY OK-KINGFISHER OK-LINCOLN OK-LOGAN OK- 
NOBLE OK-PAYNE OK- 
1118 AM CDT WED APR 22 1998 

.THIS AFTERNOON...PARTLY CLOUDY. HIGH IN THE LOWER 70S. LIGHT NORTHWEST 
WIND. 
.TONIGHT...MOSTLY CLEAR. LOW IN THE LOWER 40S. LIGHT AND VARIABLE 
WIND. 
.THURSDAY...MOSTLY SUNNY. HIGH IN THE MID 70S. LIGHT AND VARIABLE 
WIND BECOMING SOUTH IN THE AFTERNOON AND INCREASING TO 10 TO 20 MPH. 
.THURSDAY NIGHT...PARTLY CLOUDY. LOW IN THE MID 50S. 

.EXTENDED FORECAST... 
.FRIDAY...PARTLY CLOUDY AND WINDY WITH A SLIGHT CHANCE OF SHOWERS 
AND THUNDERSTORMS. HIGH IN THE UPPER 70S. 
.SATURDAY AND SUNDAY...MOSTLY CLOUDY AND WINDY WITH A CHANCE OF 
SHOWERS AND THUNDERSTORMS. LOW IN THE LOWER 50S AND HIGH IN THE 
LOWER 80S.

19980421

Dan Nelson

No BORCAL.  Partly cloudy.  No available personnel to perform calibration.

19980420

Dan Nelson

No BORCAL.  Solid cloud cover.  No available personnel to perform calibration.

19980417

Dan Nelson

No BORCAL.  Moderate Cirrus causing unstable solar conditions. Additionally, based on the attached IWIN forecast, plans for weekend BORCAL have been scrapped. 

GARFIELD OK-GRANT OK-KAY OK-KINGFISHER OK-LOGAN OK-NOBLE OK-PAYNE OK- 
430 AM CDT FRI APR 17 1998 

.TODAY...PARTLY CLOUDY MORNING...THICKER CLOUDS THIS AFTERNOON. HIGH 
IN THE UPPER 50S. NORTHEAST WIND 10 TO 15 MPH. 
.TONIGHT...MOSTLY CLOUDY WITH A 20 PERCENT CHANCE OF RAIN. LOW IN 
THE UPPER 30S. NORTHEAST WIND 5 TO 10 MPH. 
.SATURDAY...MOSTLY CLOUDY WITH A 20 PERCENT CHANCE OF RAIN. HIGH IN 
THE MID 50S. LIGHT NORTHEAST WIND. 
.SATURDAY NIGHT...MOSTLY CLOUDY. LOW IN THE LOWER 40S. 

.EXTENDED FORECAST... 
.SUNDAY...DRY. HIGH IN THE 60S. 
.MONDAY AND TUESDAY...DRY. LOW IN THE 40S AND HIGH 65 TO 75.

19980416

Dan Nelson

No BORCAL.  Broken clouds with gusting winds around 8 m/s.

19980415

Dan Nelson

No BORCAL.  Broken clouds with gusting winds around 10 m/s.

19980414

Dan Nelson

Continue preparations for BORCAL SGP 1998-01.  Still working through several RCF technical issues with Steve Wilcox at NREL. 

The last issues were taken care of at about 1900 CDST, so all that we need now are three clear sky, low wind (<5 m/s) days.

19980413

Dan Nelson

Continue preparations for BORCAL SGP 1998-01.  Still working through several RCF technical issues with Steve Wilcox at NREL.  Installed and tested T/Rh sensor received from NREL.

19980410

Dan Nelson

Continue preparations for BORCAL SGP 1998-01.  Received T/Rh sensor from NREL and will install on 19980413.  Working through several RCF technical issues (with Steve Wilcox at NREL) discovered after but not related to the installation of the T/Rh sensor.

19980409

Dan Nelson

Continue preparations for BORCAL SGP 1998-01.  Currently awaiting receipt of T/Rh sensor and clear skies to begin the BORCAL.

19980408

Dan Nelson

Continue preparations for BORCAL SGP 1998-01. Performing shade/unshade testing to confirm radiometer positioning.

19980407

Dan Nelson

Continue preparations for BORCAL SGP 1998-01.

19980406

Dan Nelson

Began installing radiometers and ventilators in preparation for BORCAL SGP 1998-01. This process will require several days to complete.

19980320

Dan Nelson

Nothing new to report.

 19980319

Dan Nelson

Preparations continue for BORCAL SGP 1998-01 to begin on or about 19980406. 

New radiometer hail shields are in place and will allow a single person to quickly cover radiometers.

Improved electrical grounding for the radiometer tables is expected to be completed by the end of next week.

The RCC system and it's associated components and software have been calibrated and tested by NREL for the upcoming calibration season. This activity was completed on 19980312 by Tom Stoffel, Steve Wilcox, Jim Treadwell, and Bev Kay.

It is currently snowing.

Lamont, OK